# Studies of Cherenkov Photon Production in PbF2 Crystals using Proton Beams at Fermilab

### v1

@@ 1 Introduction
The physics programs of future lepton colliders can be enhanced through a calorimeter system that offers state-of-the-art energy resolutions for both hadrons and photons/electrons (“EM” objects) [1, 2, 3, 4]. Calorimetric measurements of hadron energies have a resolution component arising from large event-by-event fluctuations in the fraction of “invisible energy” due to nuclear binding energies, which often do not produce detectable signals. There are three main thrusts towards improving hadronic resolution: high granularity calorimetry with particle flow (which aims to utilize the calorimeter only for neutral hadrons, while relying on tracking detectors for charged particle momentum measurement) [5], compensating hadronic calorimetry (which enhances processes such as neutron interactions correlated with nuclear breakup but which are only effective for relatively low sampling fractions) [6], and dual readout (DR) calorimetry (which uses proxies to estimate the “invisible energy” in hadron showers) [7]. The first two methods, along with the most common implementation of DR that uses plastic or quartz fibers embedded in an absorber to produce and collect signals from scintillation and Cherenkov radiation separately, are based on sampling calorimeters. These approaches suffer from a reduced resolution for EM objects, limited to approximately $13\%/\sqrt{E}$ [7].
@@
Pioneering work on DR crystal EM calorimetry was done by the DREAM collaboration [10, 11]. However, at that time, it was not possible to get sufficient Cherenkov photons without sacrificing the precision EM resolution provided by the high scintillation light output of the crystals used. In particular, their work assumed one photodetector per crystal and used the arrival time of the photons to distinguish Cherenkov photons from scintillation photons. To obtain a clear early peak from the Cherenkov photons, they had to suppress the scintillation light to a degree that photostatistics dominated its measurement, leading to EM energy resolutions comparable to those of sampling calorimeters.
@@ 2 Experimental setup
SiPMs: SiPMs (S14160-6050HS [15]) from the Hamamatsu Corporation were used to read out the raw signals from the crystal. Each SiPM module features an effective photosensitive area of $6\times 6$ mm2, with a micro-cell pitch of 50 $\mu$m and a total of 14,331 cells per module, resulting in a geometrical fill factor of 74%. These modules are of surface mount type and include a silicone window with a refractive index of 1.57. At 25C, the SiPMs have a photon detection efficiency of 50% at 450 nm [15]. However, this efficiency decreases to around 25% at 300 nm and further decreases for wavelengths below 300 nm. Similarly, the efficiency decreases from 50% to 10% as the wavelength increases from 600 nm to 800 nm. The SiPMs were powered with an over-voltage of 2.7 V.
@@
Actual setup: A photograph depicting the actual experimental setup along the beamline is shown in Fig. 4. In Fig. 4(a), the rotating platform, crystal, and readout boards are visible. To secure the crystal and boards, a 60 mm optical cage system from Thorlabs, comprising four rigid steel rods and two cage plates, was used. In addition, two plastic holders were positioned in the center to prevent the crystal from dislodging. This entire assembly was enclosed within a steel pot, serving as both a black box and a Faraday cage to shield against electromagnetic pulses, as illustrated in Fig. 4(b). A thick two-layer curtain was employed to envelop the entire setup, preventing outside light from entering. The two scintillation tiles were mounted on the pot’s surface to provide the trigger signal, and both the scintillation tiles and the crystal were aligned with the beamline.
@@
((a))
((b))
Figure 4: Actual experimental setup at Fermilab. (a) The setup with the rotating platform, crystal, and readout boards; and (b) The entire assembly was enclosed within a steel pot with two scintillation tiles mounted on the pot’s surface.
@@ 4 Results
Due to the geometric symmetry in our experimental setup, channels 0 and 7, 1 and 6, 2 and 5, and 3 and 4, are mirror images of each other. Consequently, we anticipate similar responses when the rotation angle is flipped. Figure 7 illustrates the average number of detected photons as a function of the rotation angle for channels 0 and 7. Data acquired with negative rotation angles are mirrored to positive rotation angles. It is evident that the data obtained with $-\theta$ for channel 0 closely resemble the data obtained with $+\theta$ for channel 7. Similarly, the data obtained with $-\theta$ for channel 7 closely resemble the data obtained with $+\theta$ for channel 0.
Figure 7: Average timing spectrum for channels 0 (A side) and 7 (B side). For the $x$ axis, the absolute value of the rotation angle is shown.
@@
Figure 10 shows the most probable value of the fitted peak amplitude distribution (indicating the average number of detected photons) as a function of the rotation angle $\theta$ for four channels (0, 1, 4, and 5). The data are compared to the simulation results, where the number from the simulation is normalized to the data value at $\theta=0^{\circ}$. In general, the number of detected photons increases with increasing $|\theta|$. However, this trend reverses at $\theta\simeq 70^{\circ}$, corresponding to the proton’s path length reaching its maximum at $\theta=67.4^{\circ}$.
@@
An intriguing observation is the reduction in the number of detected photons when the rotation angle is near $20^{\circ}$. Detailed simulations are conducted for $\theta=20^{\circ}$ and $\theta=10^{\circ}$ to investigate this phenomenon. Figure 11 illustrates the photon reflection angles on different surfaces for the case with $\theta=20^{\circ}$. According to geometrical optics, the incident angle of photons striking the long surface is concentrated around $36^{\circ}$ and $76^{\circ}$, as confirmed by the simulation in Fig. 12(a). Subsequently, in Fig. 12(b), the same angles are shown for photons reaching Side B of the crystal and detected by the corresponding SiPMs. These photons have an incident angle of around $14^{\circ}$ or $54^{\circ}$ on the existing surfaces. Given that the total internal reflection angle from the crystal to the silicone rubber is $51.8^{\circ}$, many of the photons with an incident angle close to $54^{\circ}$ will be reflected and thus not collected by the SiPMs. This explains the observed decrease in Cherenkov photons when $\theta$ is close to $20^{\circ}$.
For the case with $\theta=10^{\circ}$, Fig. 13 depicts the corresponding incident angles on the long surface for these photons with their first reflections and detection by the SiPMs. The majority of these photons have an incident angle of around $46^{\circ}$ or $66^{\circ}$ on the long surface and an incident angle of around $24^{\circ}$ or $44^{\circ}$ on the existing surfaces. These photons do not undergo internal reflection between the surfaces of the crystal and the silicone rubber and are detected by the SiPMs.
@@
Figure 10: Mean number of photons from $\rm PbF_{2}$ for channel 0 (a), channel 1 (b), channel 4 (c), and channel 5 (d). Each bottom panel shows the ratio of the mean number of photons measured in data and in the simulation.
Figure 11: Photon reflection angles for the case with $\theta=20^{\circ}$.
@@
It is also interesting that the asymmetry is close to 0 when $\theta=\pm 90^{\circ}$, indicating that the numbers of Cherenkov photons detected at the opposite ends of the crystal are similar. Detailed simulation studies are conducted for the case with $+90^{\circ}$. Figure 15 illustrates the photon incident angle distributions for all photons reaching side A (top left plot of Fig. 15) and those detected by the SiPMs on side A (bottom left plot of Fig. 15). The figure also shows the corresponding distributions for all photons reaching side B (top right plot of Fig. 15) and those detected by the SiPMs on side B (bottom right plot of Fig. 15). It can be observed that most generated photons reaching the two sides have incident angles concentrated around $56^{\circ}$, which is close to the internal reflective angle of $51.8^{\circ}$. Consequently, most photons will be reflected and not detected by the SiPMs. The number of detected by the SiPMs on side A is similar to the number detected by the SiPMs on side B; as a result, the asymmetry is close to 0 when $\theta=\pm 90^{\circ}$.
Figure 15: Top left: incident angles of photons hitting side A; bottom left: incident angles of photons detected by the SiPMs on side A; bottom left: incident angles of photons hitting side B; bottom right: incident angles of photons detected by the SiPMs on side B. The crystal rotation angle is $+90^{\circ}$.
@@ 5 Conclusions
A study of the Cherenkov light yield for a $\rm PbF_{2}$ crystal excited by protons as a function of the proton’s incident angle is presented. A complicated angular dependence is observed, and can be understood by considering the crystal’s geometry, contributions of index matching, total internal reflection, and non-specular reflections of the crystal, the silicon rubber, and the wrapper material. Attention to these aspects will be crucial for realizing the CalVision program. The setup we have will be useful for future DR studies using $\rm PbWO_{4}$ and BGO crystals.

### v2

@@ 1 Introduction
The physics programs of future lepton colliders can be enhanced through a calorimeter system that offers state-of-the-art energy resolutions for both hadrons and photons/electrons (“EM” objects) [1, 2, 3, 4]. Calorimetric measurements of hadron energies have a resolution component arising from large event-by-event fluctuations in the fraction of “invisible energy” due to nuclear binding energies, which often do not produce detectable signals. There are three main thrusts towards improving hadronic resolution: high granularity calorimetry with particle flow (which aims to utilize the calorimeter only for neutral hadrons, while relying on tracking detectors for charged particle momentum measurement) [5], compensating hadronic calorimetry (which enhances processes such as neutron interactions correlated with nuclear breakup but which are only effective for relatively low sampling fractions) [6], and dual readout (DR) calorimetry (which uses proxies to estimate the “invisible energy” in hadron showers) [7]. The first two methods, along with the most common implementation of DR that uses plastic and quartz (or clear) fibers embedded in an absorber to produce and collect signals from scintillation and Cherenkov radiation separately, are based on sampling calorimeters. These approaches suffer from a reduced resolution for EM objects, limited to approximately $13\%/\sqrt{E}$ [7].
@@
Pioneering work on DR crystal EM calorimetry was done by the DREAM/RD52 collaboration [10, 7]. However, at that time, it was not possible to get sufficient Cherenkov photons without sacrificing the precision EM resolution provided by the high scintillation light output of the crystals used. In particular, their work assumed one photodetector per crystal and used different optical filters and the arrival time of the photons to distinguish Cherenkov photons from scintillation photons. To obtain a clear early peak from the Cherenkov photons, they had to suppress the scintillation light to a degree that photostatistics dominated its measurement, leading to EM energy resolutions comparable to those of sampling calorimeters.
@@ 2 Experimental setup
SiPMs: SiPMs (S14160-6050HS [14]) from the Hamamatsu Corporation were used to read out the raw signals from the crystal. Each SiPM module features an effective photosensitive area of $6\times 6$ mm2, with a micro-cell pitch of 50 $\mu$m and a total of 14,331 cells per module, resulting in a geometrical fill factor of 74%. These modules are of surface mount type and include a silicone window with a refractive index of 1.57. At 25C, the SiPMs have a photon detection efficiency of 50% at 450 nm [14]. However, this efficiency decreases to around 25% at 300 nm and further decreases for wavelengths below 300 nm. Similarly, the efficiency decreases from 50% to 10% as the wavelength increases from 600 nm to 800 nm. The SiPMs were powered with an overvoltage of 2.7 V.
@@
The SiPM response was calibrated on the bench using a highly attenuated laser pulse to measure the distribution of the first several photoelectron (PE) peaks and to determine ADC-to-PE conversion. An additional gain stage (10$\sim$ 10$\times$ 10 ×) was included to ensure a 1 PE signal $>25$ ADC counts for this measurement. After correcting for the low gain configuration used in these studies the response of 0.7 mV/PE corresponded to approximately 3 ADC counts/PE.
@@
Actual setup: A photograph depicting the actual experimental setup along the beamline is shown in Fig. 4. To secure the crystal and boards, a 60 mm optical cage system from Thorlabs, comprising four rigid steel rods and two cage plates, was used. In addition, two plastic holders were positioned in the center to prevent the crystal from dislodging. This entire assembly was enclosed within a steel pot, serving as both a black box and a Faraday cage to shield against electromagnetic pulses. A thick two-layer curtain was employed to envelop the entire setup, preventing outside light from entering. The two scintillation tiles were mounted on the pot’s surface to provide the trigger signal, and both the scintillation tiles and the crystal were aligned with the beamline.
@@
Figure 4: Actual experimental setup at Fermilab. The entire assembly was enclosed within a steel pot with two scintillation tiles mounted on the pot’s surface.
@@ 4 Results
Due to the geometric symmetry in our experimental setup, channels 0 and 7, 1 and 6, 2 and 5, and 3 and 4, are mirror images of each other. Consequently, we anticipate similar responses when the rotation angle is flipped. Figure 7 illustrates the average ADC value as a function of the rotation angle for channels 0 and 7. Data acquired with negative rotation angles are mirrored to positive rotation angles. It is evident that the data obtained with $-\theta$ for channel 0 closely resemble the data obtained with $+\theta$ for channel 7. Similarly, the data obtained with $-\theta$ for channel 7 closely resemble the data obtained with $+\theta$ for channel 0.
Figure 7: Average ADC value for channels 0 (A side) and 7 (B side). Data acquired with negative rotation angles are mirrored to positive rotation angles. Points with the label “A side (+)” mean data taken with positive rotation angles ($+\theta$) for channel 0, and points with the label “A side (-)” mean data taken with negative rotation angles ($-\theta$) for channel 0. For the $x$ axis, the absolute value of the rotation angle is shown.
@@
Figure 10 shows the most probable value of the fitted peak amplitude distribution (indicating the average number of detected photons) as a function of the rotation angle $\theta$ for four channels (0, 1, 4, and 5). The data are compared to the simulation results, where the average ADC values measured in the data are converted into the numbers of photons detected. The simulation reproduces the general trends observed in the data as a function of the rotation angles for different SiPMs. However, the absolute number of photons detected can differ between the data and simulation by 30% for certain rotation angles. In general, the number of detected photons increases with increasing $|\theta|$. However, this trend reverses at $\theta\simeq 70^{\circ}$, corresponding to the proton’s path length reaching its maximum at $\theta=67.4^{\circ}$.
@@
An intriguing observation is the reduction in the number of detected photons when the rotation angle is near $20^{\circ}$. Detailed simulations are conducted for $\theta=20^{\circ}$ and $\theta=10^{\circ}$ to investigate this phenomenon. For the case with $\theta=20^{\circ}$, according to geometrical optics, the incident angle of photons striking the long surface is concentrated around $36^{\circ}$ and $76^{\circ}$, as confirmed by the simulation in Fig. 11(a). Subsequently, in Fig. 11(b), the same angles are shown for photons reaching side B of the crystal and detected by the corresponding SiPMs. These photons have an incident angle of around $14^{\circ}$ or $54^{\circ}$ on the exiting surfaces. Given that the total internal reflection angle from the crystal to the silicone rubber is $51.8^{\circ}$, many of the photons with an incident angle close to $54^{\circ}$ will be reflected and thus not collected by the SiPMs. This explains the observed decrease in Cherenkov photons when $\theta$ is close to $20^{\circ}$.
For the case with $\theta=10^{\circ}$, Fig. 12 depicts the corresponding incident angles on the long surface for these photons with their first reflections and detection by the SiPMs. The majority of these photons have an incident angle of around $46^{\circ}$ or $66^{\circ}$ on the long surface and an incident angle of around $24^{\circ}$ or $44^{\circ}$ on the exiting surfaces. These photons do not undergo internal reflection between the surfaces of the crystal and the silicone rubber and are detected by the SiPMs.
@@
Figure 10: Comparison between data and simulation for the mean number of photons detected for channel 0 (a), channel 1 (b), channel 4 (c), and channel 5 (d). Each bottom panel shows the ratio of the mean number of photons measured in data and in the simulation.
@@
It is also interesting that the asymmetry is close to 0 when $\theta=\pm 90^{\circ}$, indicating that the numbers of Cherenkov photons detected at the opposite ends of the crystal are similar. Detailed simulation studies are conducted for the case with $+90^{\circ}$. Figure 14 illustrates the photon incident angle distributions for all photons reaching side A (top left plot) and those detected by the SiPMs on side A (bottom left plot). The figure also shows the corresponding distributions for all photons reaching side B (top right plot) and those detected by the SiPMs on side B (bottom right plot). It can be observed that most generated photons reaching the two sides have incident angles concentrated around $56^{\circ}$, which is close to the internal reflective angle of $51.8^{\circ}$. Consequently, most photons will be reflected and not detected by the SiPMs. The number of photoelectrons detected by the SiPMs on side A is similar to the number of photoelectrons detected by the SiPMs on side B; as a result, the asymmetry is close to 0 when $\theta=\pm 90^{\circ}$.
Figure 14: Top left: incident angles of photons hitting side A; bottom left: incident angles of photons detected by the SiPMs on side A; top right: incident angles of photons hitting side B; bottom right: incident angles of photons detected by the SiPMs on side B. The crystal rotation angle is $+90^{\circ}$.
@@ 5 Conclusions
A study of the Cherenkov light yield for a $\rm PbF_{2}$ crystal excited by protons as a function of the proton’s incident angle is presented. A non-trivial angular dependence is observed in data for all SiPMs, and the general trend can be understood by considering the crystal’s geometry, contributions of index matching, total internal reflection, and non-specular reflections of the crystal, the silicon rubber, and the wrapper material. The simulation can reproduce data for the number of photons detected within 30%. Attention to these aspects will be crucial for realizing the CalVision program [19]. The setup we have will be useful for future DR studies using $\rm PbWO_{4}$ and BGO crystals.